If $10,000 and $4000 are invested such that $1250 is earned in one year, and if the rate of interest on the larger investment is 2% more than that of the smaller investment, find the rates of interest. The answers are $10,000 at 9.5% and $4000 at 7.5%, I just don't know how to get them.
Answer by fractalier(6550)

The setup is like this:
(10000)(x%) + 4000((x-2)%) = 125000 (we have to multiply 1250 by 100 because we are treating percents as whole numbers here)
now work it out...
10000x + 4000x - 8000 = 125000
14000x = 133000
x = 9.5%
x - 2 = 7.5%